OpenStudy (anonymous):

factor by grouping...x^3-9x^2-3x-27

OpenStudy (anonymous):

i got this far. (x^3-9x^2)-(3x-27)

OpenStudy (anonymous):

Are one of those two minus signs in front of either 3 or 27 supposed to be a plus sign? Otherwise, you can't solve this by "grouping"

OpenStudy (anonymous):

oh yes im sorry. its x^3+9x^2. maybe thats why im having trouble lol i wrote it down wrong

OpenStudy (anonymous):

Mhm, so then the answer would be \[x^3+9x^2-3x-27 = x^2(x+9)-3(x+9)=(x^2-3)(x+9)=(x-\sqrt{3})(x+\sqrt{3})(x+9)\]

OpenStudy (mertsj):